OpenStudy (anonymous):

Write the equation of the line that contains the point (-5,4) and is parallel to the line y − 3x = 2?? how do i solve?

OpenStudy (aroub):

Write down all the givens and the equation that we will be using ( and don't forget parallel means same slope) : y=mx+b; y=4, x=-5, m=-3 4=-3(-4)+b 4=12+b -8=b => y=-3x-8

OpenStudy (anonymous):

thank you

OpenStudy (aroub):

No problemm! =)

OpenStudy (anonymous):

you put -4 when x is -5 but i understand

OpenStudy (aroub):

Oh right! That's a mistake. Here's the correct solution: 4=-3(-5)+b 4=15+b -11=b => y=-3x-11